# SANSA: Unleashing the Hidden Semantics in SAM2 for Few-Shot Segmentation

### v1

@@ 4.3 Ablation Studies
In Tab. 5, we present ablation studies on COCO-20i and LVIS-92i. The first row highlights that employing SAM2 in zero-shot leads to suboptimal results. Introducing parameter-efficient tuning within the backbone unlocks SAM2 latent semantics (+24.8% on COCO, +28.2% on LVIS), while our loss, by enforcing multi-frame consistence, provides a further gain of +3.1%, and +4.1% respectively.
Ablation study COCO-20i LVIS-92i SAM2 32.2 16.5 + AdaptFormer 57.0 44.7 + Our Loss 60.1 49.8
Table 3: Effect of our contributions w.r.t. zero shot SAM2.
Adaptation Methods Params mIoU LoRa 18 M 58.0 Adapter 20 M 59.4 AdaptFormer 10 M 60.1
Table 4: Invariance to the choice of adaptation strategy.
Adapter Dimension Params mIoU 768 26 M 58.2 576 16 M 59.6 384 10 M 60.1
Table 5: Sensitivity to adapter bottleneck dimensionality.
The core strength of our method lies in the intuition that SAM2 inherently encode a latent but rich semantic structure. As a result, SANSA robust performance is not tied to the adaptation method. We experiment in Tab. 5 with the popular methods of Adapter [26], LoRa [27], and AdaptFormer [11], observing minimal performance variability. The only hyperparameter is the adapter bottleneck dimension, which we study in Tab. 5. An insight from this ablation is that lower dimensionality leads to better generalization, as it forces the adapter to focus on core semantic concepts.
Generalization Across Domain and Style Shifts. To explore how well SANSA generalizes beyond the scope of standard benchmarks, we follow [56, 37] and present qualitative examples drawn from in-the-wild image pairs, shown in Fig. 7. Our method preserves SAM2 generalization capability by learning a semantic mapping within the frozen feature space, enabling reliable correspondence even under significant distribution shifts.
@@ 5 Conclusions
In this work, we introduced SANSA, which enhances SAM2 to accept and propagate a visual prompt across frames, and shows that its memory attention mechanism can be re-focused towards semantic correspondences by restructuring the feature space. We addressed three fundamental questions: (1) we demonstrated that semantic information can be extracted from SAM2 features through lightweight bottleneck transformations, answering our initial research question, (2) by keeping SAM2 frozen, we preserve its segmentation capabilities, and finally (3) our experiments show that the learned semantic mapping generalizes robustly to novel classes, with SOTA performance on strict few-shot benchmarks and against generalist models. Beyond technical contributions, SANSA offers practical advantages: it supports diverse prompts for annotation-efficient applications, and offers a 3x speed increase. Finally, our results suggest that foundation models like SAM2 may contain richer task-adaptable knowledge than their original objectives imply, a direction worthy of future exploration.
Appendix A Appendix
@@ A.1 Limitations and Future Works
By keeping the Segment Anything 2 (SAM2) [48] model entirely frozen and storing only adapter weights, our SANSA preserves SAM2 state-of-the-art performance on Video Object Segmentation (VOS), while enabling few-shot segmentation within a unified pipeline. As our method builds directly upon SAM2, it also inherits its design assumptions and limitations. In particular, while SAM2 supports tracking and segmenting multiple objects across a video, it does so by processing each object independently. This behavior is standard in few-shot segmentation, where each object is typically queried and segmented in isolation using a separate prompt. Nevertheless, future work could explore leveraging contextual relationships between multiple objects, which may improve segmentation accuracy through cross-instance interactions and enable multi-class few-shot segmentation.
@@ A.2 Societal Impacts
Our method leverages the publicly available, pre-trained weights of Segment Anything 2 (SAM2), thus requiring minimal additional computational resources. This significantly limits the environmental impact commonly associated with training large foundation models from scratch. By providing an analysis of the underlying mechanisms of foundation models like SAM2, we contribute to a deeper understanding of their operation and potential applications beyond conventional tasks. We believe this knowledge dissemination can foster broader awareness and acceptance of foundation models both within and outside the scientific community.
@@ B.1 Analyzing Semantic Structure via Principal Component Decomposition
Our hypothesis is that SAM2 features do encode semantic information, albeit entangled with low-level signals, whereas our adapted SANSA features disentangle and make this semantic content explicit. To validate this hypothesis, we analyze their feature spaces via a clustering-based experiment, reported in Fig. 8. Specifically, we evaluate how well semantic clusters form in lower-dimensional subspaces obtained via PCA, comparing frozen SAM2 to our adapted SANSA features. To this end, we extract object-level features from fold 0 of COCO-20i, which contains categories unseen by SANSA (trained on the remaining folds). We then analyze the semantic organization of these features through a clustering-based assignment task combined with Principal Component Analysis (PCA):
@@ B.2 Semantic Discriminability and Downstream Transferability
The previous analysis revealed that in the adapted SANSA feature space, semantic information is predominantly encoded. This raises a natural follow-up question: to what extent is semantic information already encoded in the original, frozen SAM2 features, and how easily can it be accessed for downstream tasks?
To investigate this, we present a two-fold analysis in Fig. 9, comparing two models: i) frozen SAM2, and ii) our SANSA trained on folds $1,2,3$ of COCO-20i. We conduct two experiments on fold 00, i.e. on unseen classes for our model:
@@
    Linear probing: following standard practice in representation learning [19, 45, 9], we conduct a linear probing experiment to assess semantic discriminability in the two feature spaces. We extract features from both models and train a pixel-level linear classifier on the training set of COCO-20i fold 0. Note that both models are frozen and only a single linear layer is trained for the task of semantic segmentation. Evaluation is performed on the corresponding validation set. A high mean Intersection-over-Union (mIoU) indicates that semantic signals are present in the feature space and can be recovered by a simple linear projection.
@@
Figure 9: Left: Linear probing for Semantic Segmentation on a fold of COCO-20i, comparing zero-shot SAM2 and SANSA (trained on a disjoint class set). Right: Few-shot segmentation performance on the same set of data. Despite poor zero-shot performance on FSS, a simple linear layer (cf linear probing) can learn to discriminate semantic-classes from SAM2 frozen features. These results, together, suggest the presence of semantic structure, albeit entangled with other signals.
@@ B.3 Extracting Semantics Without Compromising SAM2 Capabilities
In the previous sections, we investigate how SANSA exposes the latent semantic structure embedded in frozen SAM2 features, which was our main goal. Moreover, an important principle of our design was to extract this semantic structure without sacrificing SAM2 Promptable Segmentation capabilities or its state-of-the-art performance on Video Object Segmentation. Consequently, our choice was to insert AdaptFormer blocks [11] into the last two layers of the frozen Image Encoder, which introduce residual bottleneck projections, reorganizing the feature space without altering any of the SAM2 weights.
For comparison, we evaluate several alternative adaptation strategies: i) fine-tuning only the decoder, ii) fine-tuning the Query, Key, and Value (QKV) projections within the attention layers [58], and iii) full backbone fine-tuning. Decoder fine-tuning updates around 4 million parameters, QKV fine-tuning modifies approximately 50 million parameters, and full backbone fine-tuning involves over 210 million parameters.
Tab. 6 reports results under two settings: out-of-domain, where models are trained on a subset of COCO-20i folds and evaluated on unseen classes (i.e., the strict few-shot segmentation setting, which is our primary focus), and in-domain, where models are trained on all categories.
Table 6: Comparison of adaptation strategies. We compare various strategies for adapting SAM2 in terms of performance on unseen (out-of-domain) and seen classes (in-domain). SANSA achieves the best out-of-domain mIoU with minimal parameter overhead, while preserving the SAM2 weights.
Adaptation Strategy COCO-20i Trainable SAM2 Weights in-domain out-of-domain Parameters Update Decoder Finetuning 63.7 52.1 4 M QKV-only Finetuning 73.5 55.3 50 M Backbone Finetuning 74.8 55.2 210 M AdaptFormer (SANSA) 72.0 60.1 10 M
@@
On the other hand, adapters restrict updates to a low-dimensional bottleneck [27, 26], forcing the model to prioritize generalizable transformations over spurious in-domain correlations [21], demonstrating more consistent and stable performance across both seen and unseen classes. Finally, by maintaining the SAM2 backbone and decoder entirely frozen, we ensure that the original model core capabilities, such as Promptable Segmentation and Video Object Segmentation, remain uncompromised. This is critical to deploying SANSA in practical settings, obtaining a model that fully supports both geometric and visual prompts, by storing only the adapter weights.
@@ B.4 Exploring Implicit Semantics in Traditional Video Object Segmentation Methods
Our hypothesis is that the class-agnostic training objectives of Video Object Segmentation encourage the learning of semantically meaningful representations, even without category supervision. While our main analysis focuses on SAM2, which benefits from its large-scale pretraining, we investigate whether similar semantic cues arise in traditional VOS trackers trained at smaller scale.
To this end, we evaluate three representative VOS methods, AOT [70], DeAOT [71], and MAVOS [54], alongside SAM2, on the COCO-20i 1-shot segmentation benchmark. As shown in Tab. 7, all models exhibit relatively weak performance when directly transferred to few-shot segmentation. We then apply our adaptation to each of these models, and evaluate them in the strict FSS setting on COCO-20i, i.e. focusing on their ability to generalize to previously unseen classes.
The results show a consistent behavior: despite their different training scales and capacities, all VOS models demonstrate a large performance gap between their frozen and adapted versions. For instance, MAVOS improves from 27.1% to 52.3% mIoU after adaptation, while SAM2 with Hiera-B improves from 28.0 % to 55.4%. The gap between these methods and SAM2-B is attributable to its broader generalization capabilities, stemming from its diverse and large-scale pretraining.
@@
Method Backbone #Params Zero-shot +SANSA
AOT [70] Swin-B 70 M 25.0 48.5
DeAOT [71] Swin-B 75 M 26.9 51.0
MAVOS [54] Swin-B 96 M 27.1 52.3
SAM2 [48] Hiera-B 86 M 28.0 55.4
SAM2 [48] Hiera-L 234 M 32.2 60.1
@@ Appendix C Scalable and Fast Annotation
To rigorously evaluate this capability, we simulate a large-scale annotation task on both the COCO-20i and LVIS-92i validation sets, mimicking a realistic deployment scenario in which new concepts must be annotated on the fly. For both datasets, we designate fold 0 as the test split, containing categories that are held out during training. Models are trained on the remaining folds. GF-SAM is evaluated in its original zero-shot configuration. On COCO-20i, we randomly sample 20 reference images, one per category in fold 0, and use them to guide the segmentation of 10,000 target images sampled from the same fold. On LVIS-92i, which contains a larger and more fine-grained label space, we similarly sample 92 reference classes from fold 0 and use these to segment 10,000 target images. In both cases, the reference and target sets are fixed across all methods to ensure a fair comparison. Moreover, for all the methods, reference image features are computed once and cached, avoiding repeated backbone forward passes at inference time.
@@
VRP-SAM [56] 54.8 34.9 38 min
SegIC [39] 55.1 46.7 77 min
GF-SAM [73] 59.7 38.4 58 min
@@
We benchmark SANSA against recent state-of-the-art few-shot segmentation methods, including GF-SAM [73], SegIC [39], and VRP-SAM [56], by measuring the total wall-clock time required to segment all 10,000 images and evaluating segmentation quality via mean Intersection-over-Union (mIoU) against ground truth masks. Results are summarized in Tab. 8.
SANSA achieves the best trade-off between speed and accuracy across both COCO and LVIS. In particular, it completes annotation 2–5$\times$ faster than competing approaches while maintaining the highest mIoU. On LVIS, where the label space is significantly more complex, the gains are more pronounced: SANSA outperforms all baselines by a large margin (+8.7 mIoU over the best competitor), highlighting its strong generalization ability and annotation efficiency.
Appendix D Qualitative Comparison with Prior Methods

### v2

@@ 4.4 Ablation studies
Table 3: Ablation on COCO-20i. Left: fine-tuning underperforms adaptation. Middle: simple adapters yield strong gains, while higher capacity (larger bottlenecks or added complexity, e.g., MONA) hurts generalization. Right: adapting the last two stages suffices to disentangle semantics.
Fine-tuning strategies Adaptation strategies Adapter placement
Method Params mIoU Method mIoU Stages mIoU
Frozen 0 32.2 LoRA [hu2022lora] 58.0 None 32.2
From Scratch 224 M 37.1 Adapter [houlsby2019adapter] 59.4 All 0–3 59.4
Full FT 224 M 51.6 MONA [yin20255] 56.9 Early 0–1 38.7
Decoder FT 4 M 52.1 AdaptFormer [chen2022adaptformer] Middle 1–2 57.9
QKV FT 50 M 55.3 Bottleneck $0.8\times$ channel dim 58.2 Late 2–3 60.2
Backbone FT 210 M 55.2 Bottleneck $0.5\times$ channel dim 59.6
SANSA 10 M 60.2 Bottleneck $0.3\times$ channel dim 60.2
We investigate three central aspects of our design: adaptation vs. fine-tuning, adapter architecture and capacity, and adapter placement. Results are summarized in Tab.˜3.
Why adaptation instead of fine-tuning? A central question of our work is how to distill knowledge from a pretrained model (i.e., SAM2) while preserving generalization, a key challenge in FSS. To this end, we evaluate fine-tuning strategies targeting the decoder, QKV projections, backbone, and full model, and compare them with inserting adapters into frozen weights. Results show that adaptation outperforms fine-tuning, indicating that it better preserves SAM2 pretrained priors, shifting the embedding space toward task-relevant semantics without altering the underlying representations.
What makes an adaptation strategy effective? Basic adapters such as LoRA [hu2022lora], Adapter [houlsby2019adapter], and AdaptFormer [chen2022adaptformer] all yield similar gains around $\sim$27% mIoU over frozen SAM2. The slight gains ($\sim$2%) with Adapter and AdaptFormer w.r.t. LoRA suggest that a simple non-linearity can refine this structure but not fundamentally reshape it. By contrast, increasing adapter capacity, either by enlarging the bottleneck or using more complex designs such as MONA [yin20255], reduces generalization. These results show that effective adaptation thus requires simplicity and constraint: low-capacity, bottlenecked modules best expose SAM2 latent semantics, while excessive capacity tends to overfit.
Where should adapters be placed? Prior works often insert adapters throughout the entire backbone. In our case, we find that adapting only the last two stages is sufficient, as these layers already capture high-level semantic information, which is the focus of our disentanglement objective.
Additional ablations, including backbone scale and training objective, are provided in Appendix˜D.
@@ 5 Conclusions
In this work, we introduced SANSA, which enhances SAM2 to accept and propagate a visual prompt across frames, and showed that its memory attention mechanism can be re-focused towards semantic correspondences by restructuring the feature space. We addressed three fundamental inquiries: (1) we demonstrated that semantic information can be extracted from SAM2 features through lightweight bottleneck transformations, answering our initial research question, (2) we showed that this can be achieved while keeping SAM2 frozen, thereby preserving its segmentation capabilities, and finally (3) we experimentally verified that the learned semantic mapping generalizes robustly to novel classes, with SOTA performance on strict few-shot benchmarks and against generalist models. Beyond technical contributions, SANSA offers practical advantages: it supports diverse prompts for annotation-efficient applications, and offers a 3x speed increase. Finally, our results suggest that foundation models like SAM2 may contain richer task-adaptable knowledge than their original objectives imply, a direction worthy of future exploration.
Acknowledgements. Claudia Cuttano was supported by the Sustainable Mobility Center (CNMS) which received funding from the European Union Next Generation EU (Piano Nazionale di Ripresa e Resilienza (PNRR), Missione 4 Componente 2 Investimento 1.4 "Potenziamento strutture di ricerca e creazione di "campioni nazionali di R&S" su alcune Key Enabling Technologies") with grant agreement no. CN_00000023.
Gabriele Trivigno, Giuseppe Averta and Carlo Masone were supported by FAIR - Future Artificial Intelligence Research which received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013). This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources.
Appendix
Table of Contents:
  • §A: Discussion
  • §B: Emergence of Semantic Representations
  • §C: Scalable and Fast Annotation
  • §D: Baselines and Ablation Studies
  • §E: Additional Datasets and Qualitative Comparison
  • §G: Exploratory Analysis: Negative Prompts
  • §F: Exploring Dependence on Spatial Continuity in SAM2
  • §H: Implementation and Reproducibility
Appendix A Discussion
@@ A.1 Limitations and Future Works
By keeping the Segment Anything 2 (SAM2) [ravi2024sam] model entirely frozen and storing only adapter weights, our SANSA preserves SAM2 state-of-the-art performance on Video Object Segmentation (VOS), while enabling few-shot segmentation within a unified pipeline. As our method builds directly upon SAM2, it also inherits its design assumptions and limitations. In particular, while SAM2 supports tracking and segmenting multiple objects across a video, it does so by processing each object independently. This behavior is standard in few-shot segmentation, where each object is typically queried and segmented in isolation using a separate prompt. Nevertheless, future work could explore leveraging contextual relationships between multiple objects, which may improve segmentation accuracy through cross-instance interactions and enable multi-class few-shot segmentation.
@@ A.2 Societal Impacts
Our method leverages the publicly available, pre-trained weights of Segment Anything 2 (SAM2), thus requiring minimal additional computational resources. This significantly limits the environmental impact commonly associated with training large foundation models from scratch. By providing an analysis of the underlying mechanisms of foundation models like SAM2, we contribute to a deeper understanding of their operation and potential applications beyond conventional tasks. We believe this knowledge can foster broader awareness and acceptance of foundation models both within and outside the scientific community.
@@ B.1 Analyzing Semantic Structure via Principal Component Decomposition
Our hypothesis is that SAM2 features do encode semantic information, albeit entangled with low-level signals, whereas our adapted SANSA features disentangle this semantic content to make it explicit. To validate this hypothesis, we analyze the feature spaces of both models via a clustering-based experiment, reported in Fig.˜8. Specifically, we evaluate how well semantic clusters form in lower-dimensional subspaces obtained via PCA, comparing frozen SAM2 to our adapted SANSA features. To this end, we extract object-level features from fold 0 of COCO-20i, which contains categories unseen by SANSA (trained on the remaining folds). We then analyze the semantic organization of these features through a clustering-based assignment task combined with Principal Component Analysis (PCA):
@@ B.2 Semantic Discriminability and Downstream Transferability
The previous analysis revealed that in the adapted SANSA feature space, semantic information is predominantly encoded. This raises a natural follow-up question: to what extent this semantic information was already encoded in the original, frozen SAM2 features, and how easily can it be accessed for downstream tasks?
To investigate this, we present a two-fold analysis in Fig.˜9, comparing two models: i) frozen SAM2, and ii) our SANSA trained on folds $1,2,3$ of COCO-20i. We conduct two experiments on fold 0, i.e. on unseen classes for our model:
@@
    Linear probing: following standard practice in representation learning [he2020momentum, oquab2023dinov2, caron2021emerging], we conduct a linear probing experiment to assess semantic discriminability in the two feature spaces. We extract features from both models and train a pixel-level linear classifier on the training set of COCO-20i fold 0. Note that both models are frozen and only a single linear layer is trained for the task of semantic segmentation. Evaluation is performed on the corresponding validation set. A high mean Intersection-over-Union (mIoU) indicates that semantic signals are present in the feature space and can be recovered by a simple linear projection.
@@
Figure 9: Left: Linear probing for Semantic Segmentation on a fold of COCO-20i, comparing frozen SAM2 and SANSA (trained on a disjoint class set). Right: Few-shot segmentation performance on the same set of data. Despite poor performance on few-shot segmentation, a simple linear layer (cf linear probing) can learn to discriminate semantic-classes from SAM2 frozen features. These results, together, suggest the presence of semantic structure, albeit entangled with other signals.
@@ B.3 Extracting Semantics Without Compromising SAM2 Capabilities
In the previous sections, we investigate how SANSA exposes the latent semantic structure embedded in frozen SAM2 features, which was our main goal. Moreover, an important principle of our design was to extract this semantic structure without sacrificing SAM2 Promptable Segmentation capabilities or its state-of-the-art performance on Video Object Segmentation. Consequently, our choice was to insert AdaptFormer blocks [chen2022adaptformer] into the last two layers of the frozen Image Encoder, which introduce residual bottleneck projections, reorganizing the feature space without altering any of the SAM2 weights.
For comparison, we evaluate several alternative adaptation strategies: i) fine-tuning only the decoder, ii) fine-tuning the Query, Key, and Value (QKV) projections within the attention layers [touvron2022three], and iii) full backbone fine-tuning. Decoder fine-tuning updates around 4 million parameters, QKV fine-tuning modifies approximately 50 million parameters, and full backbone fine-tuning involves over 210 million parameters.
Tab.˜4 reports results under two settings: out-of-domain, where models are trained on a subset of COCO-20i folds and evaluated on unseen classes (i.e., the strict few-shot segmentation setting, which is our primary focus), and in-domain, where models are trained on all categories.
Table 4: Comparison of adaptation strategies. We compare various strategies for adapting SAM2 in terms of generalization to unseen classes (out-of-domain) and performance on seen classes (in-domain). SANSA achieves the best out-of-domain mIoU with minimal parameter overhead, while preserving the original SAM2 weights.
Adaptation Strategy COCO-20i Trainable SAM2 Weights
in-domain out-of-domain Parameters Update
Decoder Finetuning 63.7 52.1 4 M
QKV-only Finetuning 73.5 55.3 50 M
Backbone Finetuning 74.8 55.2 210 M
AdaptFormer (SANSA) 72.0 60.2 10 M
@@
On the other hand, adapters restrict updates to a low-dimensional bottleneck [hu2022lora, houlsby2019adapter], forcing the model to prioritize generalizable transformations over spurious in-domain correlations [he2021effectiveness], demonstrating more consistent and stable performance across both seen and unseen classes.
Finally, by maintaining the SAM2 backbone and decoder entirely frozen, we ensure that the original model core capabilities, such as Promptable Segmentation and Video Object Segmentation, remain uncompromised. This is critical to deploying SANSA in practical settings, obtaining a model that fully supports both geometric and visual prompts, by storing only the adapter weights.
@@ B.4 Exploring Implicit Semantics in Traditional Video Object Segmentation Methods
Our hypothesis is that the class-agnostic training objective of Video Object Segmentation encourages the learning of semantically meaningful representations, even without category supervision. While our main analysis is focused on SAM2, which benefits from its large-scale pretraining, we investigate whether similar semantic understanding properties arise in traditional VOS trackers trained at smaller scale.
To this end, we evaluate three representative VOS methods, AOT [yang2021associating], DeAOT [yang2022decoupling], and MAVOS [shaker2024efficient], alongside SAM2, on the COCO-20i 1-shot segmentation benchmark. As shown in Tab.˜5, all models exhibit relatively weak performance when directly transferred to few-shot segmentation. We then apply our adaptation to each of these models, and evaluate them in the strict FSS setting on COCO-20i, i.e. focusing on their ability to generalize to previously unseen classes. The results show a consistent behavior: despite their different training scales and capacities, all VOS models demonstrate a large performance gap between their frozen and adapted versions. For instance, MAVOS improves from 27.1% to 52.3% mIoU after adaptation, while SAM2 with Hiera-B improves from 28.0 % to 55.4%. The gap between these methods and SAM2-B is attributable to its broader generalization capabilities, stemming from its diverse and large-scale pretraining.
@@
Method Backbone #Params Frozen +SANSA
AOT [yang2021associating] Swin-B 70 M 25.0 48.5
DeAOT [yang2022decoupling] Swin-B 75 M 26.9 51.0
MAVOS [shaker2024efficient] Swin-B 96 M 27.1 52.3
SAM2 [ravi2024sam] Hiera-B 86 M 28.0 55.4
SAM2 [ravi2024sam] Hiera-L 234 M 32.2 60.2
@@ Appendix C Scalable and Fast Annotation
To rigorously evaluate this capability, we simulate a large-scale annotation task on both the COCO-20i and LVIS-92i validation sets, mimicking a realistic deployment scenario in which new concepts must be annotated on the fly. For both datasets, we designate fold 0 as the test split, containing categories that are held out during training. Models are trained on the remaining folds. On COCO-20i, we randomly sample 20 reference images, one per category in fold 0, and use them to guide the segmentation of 10,000 target images sampled from the same fold. On LVIS-92i, which contains a larger and more fine-grained label space, we similarly sample 92 reference classes from fold 0 and use these to segment 10,000 target images. In both cases, the reference and target sets are fixed across all methods to ensure a fair comparison. Moreover, for all the methods, reference image features are computed once and cached, avoiding repeated backbone forward passes at inference time.
@@
VRP-SAM [sun2024vrp] 54.8 34.9 38 min
SegIC [meng2024segic] 55.1 46.7 77 min
GF-SAM [zhang2024bridge] 59.7 38.4 58 min
@@
We benchmark SANSA against recent state-of-the-art few-shot segmentation methods, including GF-SAM [zhang2024bridge], SegIC [meng2024segic], and VRP-SAM [sun2024vrp], by measuring the total wall-clock time required to segment all 10,000 images and evaluating segmentation quality via mean Intersection-over-Union (mIoU) against ground truth masks. Results are summarized in Tab.˜6.
[…]
Appendix D Baselines and Ablation Studies
D.1 Evaluating the Impact of SAM2 on Prior SAM-Based Pipelines
We analyze the effect of replacing SAM with SAM2 in existing Segment Anything-based methods, namely Matcher [liu2023matcher], GF-SAM [zhang2024bridge], and VRP-SAM [sun2024vrp]. These methods follow a modular two-stage design: an external backbone (e.g., DINOv2 for Matcher and GF-SAM, ResNet-50 for VRP-SAM) is used to perform feature matching, and the matched features are then used to generate geometric prompts that are passed to SAM, which operates at the image level to predict the final segmentation mask. We re-evaluate these approaches by replacing SAM with SAM2 in the final segmentation step, while leaving the matching backbone unchanged. Note that all these methods adopt the largest version of SAM (SAM-H), hence we replace it with the largest SAM2 version (SAM2-L). As shown in Tab.˜7, the performance differences are negligible. This is consistent with findings from the SAM2 paper [ravi2024sam], which shows that SAM2-L yields only little gains over SAM-H in image-level promptable segmentation tasks. In other words, given the same prompts the segmentation quality is essentially unchanged, as both SAM and SAM2 excel at converting prompt geometry into accurate masks.
Table 7: Performance comparison of Segment Anything-based methods replacing SAM with SAM2. These methods prompt SAM at image-level, hence SAM2 additional capabilities (e.g. mask propagation) play no role in this setting. As a result, SAM2 provides only minimal gains for existing two-stage methods (Matcher [liu2023matcher], GF-SAM [zhang2024bridge], VRP-SAM [sun2024vrp]).
Method Feature Matching Segmentation mIoU
Backbone Model (%)
PerSAM [zhang2023personalize] SAM 23.0
SAM2 23.6
Matcher [liu2023matcher] DINOv2 SAM 52.7
SAM2 52.4
VRP-SAM [sun2024vrp] ResNet50 SAM 53.9
SAM2 54.3
GF-SAM [zhang2024bridge] DINOv2 SAM 58.7
SAM2 58.9
SANSA (ours) SAM2 60.2
These findings highlight the fundamental difference between SANSA and previous SAM-based approaches. While prior methods rely on external backbones for matching and treat SAM as a image-level decoder, our architecture tightly integrates all components into a unified pipeline. By leveraging SAM2 not only for mask prediction but also for feature matching and prompt encoding, SANSA ensures more effective feature reuse, leading to stronger performance and improved computational efficiency.
D.2 Experiments with Different Backbones
In this section, we evaluate the performance of different SAM2 Encoder variants at varying scales. Table 8 presents a comparison of model parameters, inference speed (Frame Per Second), and mIoU across three versions of the model (Tiny, Base, and Large), with Frame Per Second (img/s) values computed on an NVIDIA RTX 4090.
Table 8: Comparison of model parameters, inference speed (FPS), and mIoU for different scales of the SAM2 encoder (Tiny, Base, Large) using Hiera. FPS are measured on an NVIDIA RTX 4090.
SANSA Total Trainable Frames Per Second mIoU
Encoder Parameters Parameters (img/s) (%)
Tiny 40 M 1.3 M 50 51.1
Base 86 M 3.3 M 31 55.4
Large 234 M 10 M 15 60.2
All variants of SANSA offer a favorable trade-off between efficiency and accuracy, with consistent gains in mIoU as model capacity increases. Notably, the Large model achieves the highest mIoU of 60.2%, while still maintaining a practical inference speed of 15 FPS.
D.3 Effect of the Training Objective
We next study the impact of our sequential conditioning objective. When $J=1$, training reduces to the standard few-shot segmentation objective, where each target is predicted only from the reference image, without propagation. In contrast, we propagate predictions across multiple unlabeled targets ($J>1$), encouraging the model to extract semantic structure that generalizes beyond individual pairs.
Table 9 shows that sequential conditioning consistently improves performance over the baseline with $J=1$. We observe the largest gains with J=3J=344, after which results plateau.
Table 9: Ablation on training sequence length $J$. Using $J>1$ improves over the standard few-shot objective ($J=1$). Gains saturate at J=3J=344, which we use as default.
$J$ COCO-20i (1-shot) LVIS-92i (1-shot)
1 (no propagation) 57.0 44.7
2 58.9 47.2
3 (ours) 60.2 49.6
4 59.8 49.6
5 59.5 49.8
7 60.0 49.4
At inference, we emphasize that sequential conditioning is not used: each target is segmented independently given the annotated reference(s), following the standard few-shot segmentation protocol.
Appendix E Additional Datasets and Qualitative Comparison
E.1 Qualitative Comparison with Prior Methods